# Radiation Pressure Induced Oscillations of an Optically Levitating Mirror

### 01 Jan 2024 v1

@@ I Introduction
The idea that cavity optomechanical phenomena can be observed in a cavity formed by optically levitating mirror has been proposed and realized by a number of authorsSingh et al. (2010); Guccione et al. (2013); Ma et al. (2020). The authors of Ref. Lecamwasam et al. (2020) derived main equations describing the motion of the center of mass of a levitated mirror and studied stability of its dynamics. They found that even if the system possess a stable equlibrium in the quasi-static approximation, the oscillations around this equilibrium become unstable if one takes into account corrections to the quasi-stationary approximation. These corrections result in optically induced amplification of mechanical motion and run-away instability. In this paper we show that the mechanical dissipation of the mirror, which is present due to the air resistance and can be controlled, stabilizes the nonlinear dynamics of the mirror resulting in a multimode phonon lasing-like behavior(Rivlin and Zadernovsky, 1992; Vahala et al., 2009; Mahboob et al., 2013; Yamaguchi et al., 2014; Okuyama et al., 2014; Loerch; Pettit et al., 2019; Sharma et al., 2022; Behrle et al., 2023; El-Sayed et al., 2023). In this note we present the results of our analysis of the periodic oscillations of the levitating mirror in this regime.
@@
where $P_{in}$ is the input power. Unlike standard optomechanical models(Carmon et al., 2005; Kippenberg and Vahala, 2007), Eq. 1 lacks a mechanical spring force, so that the oscillations of the mirror occur solely due to so-called “optical spring” effect(Aspelmeyer et al., 2014). It shall be noted that while mechanical spring is linear and instantaneous, the “optical spring” is non-linear with respect to mirror’s displacement, and is also characterized by a time delay determined by the lifetime of the optical cavity mode.

### 10 Jan 2024 v3

@@ I Introduction
The idea that cavity optomechanical phenomena can be observed in a cavity formed by optically levitating mirror has been proposed and realized by a number of authorsSingh et al. (2010); Guccione et al. (2013); Ma et al. (2020). The authors of Ref. Lecamwasam et al. (2020) derived main equations describing the motion of the center of mass of a levitated mirror and studied stability of its dynamics. They found that even if the system possess a stable equlibrium in the quasi-static approximation, the oscillations around this equilibrium become unstable if one takes into account corrections to the quasi-stationary approximation. These corrections result in optically induced amplification of mechanical motion and run-away instability. In this paper we show that the mechanical dissipation of the mirror, which is present due to the air resistance and can be controlled, stabilizes the nonlinear dynamics of the mirror resulting in a multimode phonon lasing-like behaviorRivlin and Zadernovsky (1992); Vahala et al. (2009); Mahboob et al. (2013); Yamaguchi et al. (2014); Okuyama et al. (2014); Pettit et al. (2019); Sharma et al. (2022); Behrle et al. (2023); El-Sayed et al. (2023). In this note we present the results of our analysis of the periodic oscillations of the levitating mirror in this regime.
@@
where $P_{in}$ is the input power. Unlike standard optomechanical modelsCarmon et al. (2005); Kippenberg and Vahala (2007), Eq. 1 lacks a mechanical spring force, so that the oscillations of the mirror occur solely due to so-called “optical spring” effectAspelmeyer et al. (2014). It shall be noted that while mechanical spring is linear and instantaneous, the “optical spring” is non-linear with respect to mirror’s displacement, and is also characterized by a time delay determined by the lifetime of the optical cavity mode.